# 3DMambaIPF: A State Space Model for Iterative Point Cloud Filtering via Differentiable Rendering

### v1

@@ Abstract.
Noise is an inevitable aspect of point cloud acquisition, necessitating filtering as a fundamental task within the realm of 3D vision. Existing learning-based filtering methods have shown promising capabilities on small-scale synthetic or real-world datasets. Nonetheless, the effectiveness of these methods is constrained when dealing with a substantial quantity of point clouds. This limitation primarily stems from their limited denoising capabilities for large-scale point clouds and their inclination to generate noisy outliers after denoising. The recent introduction of State Space Models (SSMs) for long sequence modeling in Natural Language Processing (NLP) presents a promising solution for handling large-scale data. Encouraged by iterative point cloud filtering methods, we introduce 3DMambaIPF, firstly incorporating Mamba (Selective SSM) architecture to sequentially handle extensive point clouds from large scenes, capitalizing on its strengths in selective input processing and long sequence modeling capabilities. Additionally, we integrate a robust and fast differentiable rendering loss to constrain the noisy points around the surface. In contrast to previous methodologies, this differentiable rendering loss enhances the visual realism of denoised geometric structures and aligns point cloud boundaries more closely with those observed in real-world objects. Extensive evaluations on datasets comprising small-scale synthetic and real-world models (typically with up to 50K points) demonstrate that our method achieves state-of-the-art results. Moreover, we showcase the superior scalability and efficiency of our method on large-scale models with about 500K points, where the majority of the existing learning-based denoising methods are unable to handle.
Point cloud filtering, denoising, state space model, differentiable rendering
1. Introduction
@@
Point clouds, assemblies of numerous three-dimensional points denoting spatial positions, play a crucial role in representing scenes across multimedia and computer vision (CV) domains. Nevertheless, as point clouds from real-world environments are captured by devices such as LiDAR, depth cameras, or 3D scanners, noises are inevitably introduced during the scanning process (Fei et al., 2024b). The presence of these noises undermines the required data consistency and accuracy in the application scenarios of point clouds, making filtering of point clouds a fundamental task within the realm of 3D vision (Fei et al., 2022, 2023).
@@
Due to the importance of point cloud filtering, numerous methods have been dedicated to improving the capabilities. Traditional point cloud filtering methods rely on expert priors, but they are limited by the accuracy of these priors and are not suitable for complex and sparse point clouds. In recent years, deep learning-based point cloud filtering methods have emerged prominently. These methods commonly employ iterative modules that consist of stacked encoder-decoder pairs. Such architectures leverage the output of the previous module as the input for the subsequent module, thereby demonstrating superior performance. For instance, ScoreDenoise (Luo and Hu, 2021) utilizes gradient ascent techniques and introduces an iterative point-moving method based on estimated scores. IterativePFN (de Silva Edirimuni et al., 2023) employs iterative modules and adaptive loss functions to iteratively and progressively refine noisy points toward adaptive ground truth (GT), yielding promising results on both synthetic and real-world datasets.
@@
However, these point cloud filtering methods are only applicable to small-scale scenes and not suitable for denoising large-scale scenes. It is found experimentally that some denoising methods exhibit inadequate filtering capabilities when applied to large-scale point clouds. Furthermore, previous works only have demonstrated satisfactory performance solely in low-noise environments, but they are limited to handling details in high-noise environments. This limitation is particularly evident in the generation of unrealistic boundaries, as well as a lack of appropriate constraints. Consequently, the main challenges are twofold: (i) denoising for large-scale point clouds or complex geometries, (ii) denoising for high-noise environments.
@@
Recently, State Space Models (SSMs), particularly structured SSMs (S4) (Gu et al., 2021a), have shown great promise in Natural Language Processing (NLP) (Lu et al., 2024). Built upon the S4 framework, Mamba (Gu and Dao, 2023) exhibits outstanding performance across many tasks in NLP and 2D CV  (He et al., 2024; Yang et al., 2024; Zhao et al., 2024; Misra et al., 2024). These achievements of Mamba can be ascribed to the linear computational complexity and its proficient long-range context learning capabilities on long sequences. Furthermore, the utilization of time-varying parameters and a hardware-aware algorithm by Mamba represents significant advancements in sequence modeling.
@@
Considering the aforementioned challenges in point cloud filtering and the strength of Mamba for long sequence modeling, we propose 3DMambaIPF, a novel Iterative Point Clouds Filtering model utilizing the Mamba module with point rendering techniques. Specifically, 3DMambaIPF is composed of multiple iterations of Mamba-Denoise Modules, each comprising a pair of Mamba encoders and decoders. Initially, the input point cloud is divided into patches, within which a graph structure is generated based on points and their neighbors to extract position features. In the Mamba-Encoder, the Dynamic EdgeConv module encodes positional features with MLPs, which are then inputted into a Mamba module to select and generate position-dependent features from sequential inputs. The Mamba-Decoder is responsible for upsampling features into point cloud patches. During the training process, a novel rendering loss is introduced leveraging a differential point rendering technique, and the loss is computed by the disparities of rendered images of filtered point cloud and GT. Subsequently, an adaptive GT altering throughout the iterations is introduced, gradually transitioning from an initially noise-added state to the true GT. Following the completion of all iterations, a patch-stitching method is employed to reconstruct the denoised point cloud from the patches. Experimentally, 3DMambaIPF is trained using PU-Net (Yu et al., 2018) dataset and achieves state-of-the-art results compared to off-the-shelf methods. 3DMabaIPF is also able to well generalize to high-noise point clouds from the Stanford 3D Scanning Repository (Curless and Levoy, 1996; Krishnamurthy and Levoy, 1996) for synthetic objects with complex geometries. Moreover, 3DMambaIPF outperforms baseline methods in terms of visual perceptions on real-world scene datasets. In summary, our specific contributions are as follows:
  • Mamba module is integrated into the iterative filtering module, enabling modeling and filtering of long-sequence point cloud features, thereby achieving more accurate and rapid modeling of large-scale point clouds.
@@
  • A novel, differentiable point rendering loss is integrated into 3DMambaIPF, supplanting the distance-based loss functions employed in previous methods. This enhancement endows 3DMambaIPF with the ability to handle edge denoising and makes the denoising results more realistic.
@@
  • 3DMambaIPF surpasses baseline methods by achieving state-of-the-art results on both synthetic and real-world scene datasets, while demonstrating excellent performance even on high-noise large-scale synthetic dataset.
2. Related Work
@@
Figure 1. Overview of 3DMambaIPF. An Encoder-Decoder-based model built on Mamba named Mamba-Denoise Module is introduced for iterative filtering. The input noisy point cloud is partitioned into patches and fed into the iterative Mamba-Denoise Module. Upon completion of the iterations, clean point cloud patches are produced as output.
Figure 2. A Gaussian noise with a variable standard deviation is added to the GT to generate the adaptive GT. As the standard deviation gradually decreases, the adaptive GT approaches the GT with each iteration. Eventually, the GT no longer changes in the final iteration.
2.1. Point Cloud Filtering Methods
Traditional point cloud filtering methods can be categorized into several main classes. Geometry-based approaches (Meek and Walton, 2000; Alexa et al., 2001, 2003; Cazals and Pouget, 2005; Huang et al., 2013; Lipman et al., 2007) involve projecting point clouds onto a surface and utilize geometric features, such as parameters derived from local surface fitting and normal estimation, to describe the shape and structure of the point cloud. Statistical-based methods (Avron et al., 2010; Sun et al., 2015; Leal et al., 2020; Zaman et al., 2017; Schall et al., 2005; Thakran and Toshniwal, 2012) employ techniques such as sparse reconstruction, density estimation and clustering analysis. Moreover, graph-based methods (Gao et al., 2018; Hu et al., 2020, 2021) utilize graph structures to represent the relationships between points. These methods leverage graph filters and graph signal processing techniques to denoise point clouds effectively. By modeling point clouds as graphs, where each point is a node and connections between points represent geometric relationships, graph-based methods can exploit the inherent connectivity and structure of the data for noise reduction. Nonetheless, traditional methods consistently prove inadequate in effectively denoising point clouds with large-scale and complex geometries.
@@ 2.1. Point Cloud Filtering Methods
Deep Learning-based point cloud filtering methods are undergoing significant development to address the limitations of traditional approaches. The well-known PCN (PointCleanNet) (Rakotosaona et al., 2020) model is based on PCPNet (Guerrero et al., 2018) and focuses on removing outliers and reducing noise in unordered point clouds. However, due to PCN’s denoising being performed point-wise rather than patch-wise, its efficiency is relatively low. PD-Flow (Mao et al., 2022) learns the distribution of noisy point sets and achieves denoising through noise disentanglement. However, the number of iterations needed for PD-Flow varies depending on the level of noise to achieve optimal results. GPDNet (Pistilli et al., 2020) employs graph-convolutional layers to dynamically build neighborhood graphs and utilizes the similarity of high-dimensional feature representations to construct intricate feature hierarchies. ScoreDenoise (Luo and Hu, 2021) leverages the distribution model and score estimation from noisy point clouds. Pointfilter (Zhang et al., 2020) employs an encoder-decoder architecture, where points and their neighbors are taken as input for point-wise learning. It generates a displacement vector to accomplish denoising while preserving sharp features. IterativePFN (de Silva Edirimuni et al., 2023) leverages iterative filtering modules and adaptive GT progressively yet swiftly converges noisy points onto clean surfaces, achieving superior performance. Current point cloud filtering methods are primarily designed for small-scale scenes (about 50K points) and struggle to effectively denoise point clouds with dense noise, particularly those far from clean surfaces, especially in large-scale scenes (about 500K points). Furthermore, existing methods (Zhang et al., 2020) demonstrate limited effectiveness in filtering the edges of point clouds and capturing fine-grained characteristics. Considering these challenges, our work aims to introduce a model tailored for large-scale scenes characterized by high-density noise and robust denoising capabilities at point cloud edges.
2.2. SSM-based Methods
@@
SSMs provide a robust mathematical framework rooted in control theory and are increasingly applied in fields such as NLP and CV to model complex temporal dynamics within dynamic systems. This novel modeling approach addresses the challenge of handling long sequences, a difficulty encountered by both Transformers and RNNs. By mitigating the quadratic computational complexity of Transformers and addressing the forgetting issue of RNNs, SSMs prove highly efficient in capturing temporal dependencies within sequential data.
One noteworthy advancement in SSMs is the introduction of the HiPPO (Gu et al., 2020), which simplifies input sequence modeling by compressing inputs into a dynamic, polynomial-based representation. However, despite this innovative approach, high computational and memory footprint limit the applicability as a general sequence modeling solution (Gu et al., 2021b), failing to fully address the bottleneck issues inherent in SSMs. To this end, the S4 (Gu et al., 2021a) presents a novel parameterization technique. By integrating a low-rank structured correction, S4 achieves stable diagonalization, significantly enhancing computational efficiency through simplified Cauchy kernel operations.
@@
Significant advancements have been made in the Mamba framework (Gu and Dao, 2023), which integrates selective SSMs into a simplified architecture. This integration allows Mamba to achieve rapid inference, linear scalability in sequence length, and state-of-the-art performance across multiple modalities, while also enhancing the selectivity of state representation. These advancements collectively demonstrate the ongoing evolution of SSMs and their increasing significance in modeling temporal dynamics across diverse domains. Methods based on Mamba have achieved state-of-the-art results in various fields, including large language models (He et al., 2024; Yang et al., 2024), medical image analysis (Liao et al., 2024; Yue and Li, 2024; Huang et al., 2024), and 3D point cloud perception (Liang et al., 2024; Zhang et al., 2024; Liu et al., 2024). Taking Consideration of the advantages of Mamba in modeling long sequences, we introduce Mamba in our work to address point cloud modeling problems in large-scale scenes.
2.3. Differentiable Rendering Techniques
@@
Rendering techniques encompass algorithms employed to transform three-dimensional models into two-dimensional images. The differentiable rendering methods are designed for introduction into deep learning, facilitating 3D reconstruction through error backpropagation. Current differentiable rendering methods can be categorized into four types based on their geometric representations: point-based (Insafutdinov and Dosovitskiy, 2018; Yifan et al., 2019; Müller et al., 2022; Kerbl et al., 2023; Fei et al., 2024a), triangle mesh-based (Laine et al., 2020; Liu et al., 2019; Hermosilla et al., 2018), implicit function-based(Bangaru et al., 2022; Vicini et al., 2022), and volume-based(Mildenhall et al., 2021; Ren et al., 2023). In this paper, we utilize an efficient differentiable point-based rendering method (Insafutdinov and Dosovitskiy, 2018) to render both the filtered and ground-truth point clouds during training to generate the view loss, which is then propagated backward to adjust the network parameters.
3. Preliminaries
Structured State Space Models are two-stage sequence-to-sequence models that map the input $x(t)\in\mathbb{R}^{L}$ to output $y(t)\in\mathbb{R}^{L}$. Specifically, the SSM process can be described as follows:
(1) $h^{\prime}(t)=\bf{A}$ $h(t)+{\bf{B}}x(t),$
$y(t)$ $={\bf{C}}h(t),$
where $h(t)\in\mathbb{R}^{N}$ represents the latent state, and $h^{\prime}(t)\in\mathbb{R}^{N}$ is the derivative of $h(t)$. ${\bf{A}}\in\mathbb{R}^{N\times{N}}$, ${\bf{B}}\in\mathbb{R}^{N\times{1}}$, ${\bf{C}}\in\mathbb{R}^{1\times{N}}$ are the parameters. The first stage of S4 discretizes the parameter ${\bf{A}}$, ${\bf{B}}$ into ${\bf{\bar{A}}}$, ${\bf{\bar{B}}}$ with a parameter $\Delta$:
(2) ${\bf{\bar{A}}}=exp$ $(\Delta{\bf{A}}),$
${\bf{\bar{B}}}=(\Delta{\bf{A}})^{-1}(exp$ $(\Delta{\bf{A}})-I)\cdot{\Delta{\bf{B}}}.$
After discretizing, the second stage is to calculate recursively:
(3) $h_{t}={\bf{\bar{A}}}$ $h_{t-1}+{\bf{\bar{B}}}x_{t},$
$y_{t}$ $={\bf{C}}h_{t},$
with the convolution formula:
(4) ${\bf{\bar{K}}}=({\bf{C}}\bf{\bar{B}},{\bf{C}}{\bf{\bar{A}}}{\bf{\bar{B}}}$ $,...,{\bf{C}}{\bf{\bar{A}}}^{k}{\bf{\bar{B}}},...),$
$y=x$$*{\bf{\bar{K}}}.$
@@ 3. Preliminaries
Figure 3. Recurrent Point Cloud Filtering of 3DMambaIPF. The Mamba Encoder comprises Dynamic EdgeConv modules and Mamba Blocks, while the Mamba Decoder consists of Mamba-Decoder Blocks and an Activated Linear module.
Mamba, as mentioned in Section 2.2, employs recurrent scans coupled with a selection mechanism to regulate the passage of specific segments within the sequence into the latent states. Specifically, by making the SSM parameters functions of the input, Mamba addresses the weakness in dealing with discrete modalities, allowing the model to selectively propagate or forget information based on the current token. Simultaneously, Mamba has devised a hardware-aware parallel algorithm to operate in recursive mode, addressing the challenge of inefficient convolution operations. The structure diagram of a single Mamba block is illustrated in Figure 4, where the overall process can be viewed as a transition from $x_{t-1}\in\mathbb{R}^{L}$ to $x_{t}\in\mathbb{R}^{L}$:
(5) ${x_{t}}^{\prime}=\text{DWConv}(\text{MLP}$ $(\text{LN}(x_{t-1}))),$
$s_{t}=\text{MLP}(\text{LN}(\text{SSM}(\sigma({x_{t}}$ )))×${}^{\prime})))\times\sigma(\text{LN}(x_{t-1}))),$
$x_{t}=s_{t}+x$ ${}_{t-1},$
where DWConv represents depth-wise convolution, LN represents layer normalization, $\sigma$ denotes activation (SiLU here).
4. Method
4.1. Overview
@@
In this section, building on top of IterativePFN (de Silva Edirimuni et al., 2023), we will detailly introduce 3DMambaIPF which is the first Mamba framework for iteratively filtering 3D point clouds via differentiable point rendering. The recurrent point cloud filtering structure of the 3DMambaIPF model is depicted in Figure 1. 3DMambaIPF network consists of a cascade of Mamba encoders and Mamba decoders, with each pair referred to as a Mamba-Denoise Module. After receiving an extracted patch of the point cloud as input, a single Mamba-Denoise Module constructs a directed graph from the points within the patch. The directed graph is encoded into token sequences to extract features. The denoised point token sequences are then decoded, eliminating noise, and serve as the input for the next Mamba-Denoise Module. Subsequently, we will employ a differentiable 3D point rendering loss to refine the positions of noisy points. This will be achieved by comparing the rendered images from denoised point clouds with the GT. This process iterates until the final module is reached, after which the final output is a clean point cloud patch. The recurrent point cloud filtering method will be introduced in Section 4.2, the Mamba-Denoising encoder and decoder module will be introduced in Section 4.3. Section 4.4 will introduce differentiable point rendering techniques and backpropagation, while Section 4.5 will specifically describe loss function used in 3DMambaIPF.
4.2. Recurrent Point Cloud Filtering
@@
Given a 3-dimensional clean point cloud, denoted as $X_{c}=\{x_{c1},x_{c2},...x_{cn}\}\subseteq\mathbb{R}^{3}$, each point in $X_{c}$ can be represented by three-dimensional coordinates $x_{ci}=(\alpha_{i},\beta_{i},\gamma_{i})$. To prepare the point cloud as input for the network, the K-nearest neighbors (KNN) algorithm (Peterson, 2009) is used to partition the point cloud into patches, and a directed graph is constructed within each patch. Considering a point cloud patch $X\subseteq\mathbb{K}^{3}$ which generated by KNN and a directed graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}=\{x_{1},x_{2},...x_{K}\}$ and $\mathcal{E}$ represents the edges between $x_{i}\in\mathcal{V}$ and its $k$-nearest neighbors.
@@
Figure 4. Detailed network architectures for Dynamic EdgeConv, Mamba Block, and Mamba Decoder Block.
4.3. Mamba-Denoising Backbone
@@
The Mamba-Denoising Backbone is the core component of 3DMambaIPF for recurrent point cloud filtering. As illustrated in Figure 3, the Mamba-Denoising backbone adopts an Encoder-Decoder structure. The encoder consists of four Dynamic EdgeConv modules and four corresponding Mamba Blocks respectively, while the decoder comprises a Mamba Decoder Block and a linear activation layer. The detailed network structure is illustrated in Figure 4.
@@
The Dynamic EdgeConv module facilitates the conversion of input patches into a graph structure, subsequently processed through an MLP network to extract pertinent features. Initially, vertices undergo feature extraction via an MLP, capturing intrinsic characteristics. Subsequently, an additional MLP focuses on extracting positional features from vertices connected by edges to the target vertex, facilitating feature differentiation. This process culminates in concatenating the target vertex’s features with the differences of features computed from all connected vertices, effectively deriving nuanced positional features within the patch. Following the feature extraction of the point cloud within the patch using MLPs, the resultant feature sequence is sent to the Mamba Encoder. This encoder selectively processes the input feature sequence and alters it into an output feature sequence. In the Decoder module, the Mamba Decoder Block selectively processes the input feature sequence to reduce dimensionality, gradually reconstructing it into three dimensions. Linear activation is used to determine an output movement distance of a noisy point towards a clean surface after filtering. Distance values are then combined with a residual connection to the input patch, yielding the denoised point cloud patch.
4.4. Differentiable Rendering and Backpropagation

### v2

@@ Abstract
Noise is an inevitable aspect of point cloud acquisition, necessitating filtering as a fundamental task within the realm of 3D vision. Existing learning-based filtering methods have shown promising capabilities on commonly used datasets. Nonetheless, the effectiveness of these methods is constrained when dealing with a substantial quantity of point clouds. This limitation primarily stems from their limited denoising capabilities for dense and large-scale point clouds and their inclination to generate noisy outliers after denoising. To deal with this challenge, we introduce 3DMambaIPF, for the first time, exploiting Selective State Space Models (SSMs) architecture to handle highly-dense and large-scale point clouds, capitalizing on its strengths in selective input processing and large context modeling capabilities. Additionally, we present a robust and fast differentiable rendering loss to constrain the noisy points around the surface. In contrast to previous methodologies, this differentiable rendering loss enhances the visual realism of denoised geometric structures and aligns point cloud boundaries more closely with those observed in real-world objects. Extensive evaluations on commonly used datasets (typically with up to 50K points) demonstrate that 3DMambaIPF achieves state-of-the-art results. Moreover, we showcase the superior scalability and efficiency of 3DMambaIPF on highly dense and large-scale point clouds with up to 500K points compared to off-the-shelf methods.
Codehttps://github.com/TsingyuanChou/3DMambaIPF.
Introduction
@@
Point clouds, assemblies of numerous three-dimensional points denoting spatial positions, play a crucial role in representing scenes across multimedia and computer vision (CV) domains. Nevertheless, as point clouds from real-world environments are captured by devices such as LiDAR, depth cameras, or 3D scanners, noises are inevitably introduced during the scanning process (Fei et al. 2024b). The presence of these noises undermines the required data consistency and accuracy in the application scenarios of point clouds, making filtering of point clouds a fundamental task within the realm of 3D vision (Fei et al. 2022, 2023). Additionally, due to the potential large-scale nature of these real-world scans, point clouds can exhibit substantial coordinate ranges and be highly dense, containing a significant number of points.
@@
Traditional point cloud filtering methods rely on expert priors, but they are limited by the accuracy of these priors and are not suitable for complex and sparse point clouds. In recent years, deep learning-based point cloud filtering methods have emerged prominently. These methods commonly employ iterative modules that consist of stacked encoder-decoder pairs. Such architectures leverage the output of the previous module as the input for the subsequent module, thereby demonstrating superior performance. For instance, IterativePFN (de Silva Edirimuni et al. 2023) employs iterative modules and adaptive loss functions to iteratively and progressively refine noisy points toward adaptive ground truth (GT), yielding promising results on both synthetic and real-world datasets.
@@
However, these point cloud filtering methods are only applicable to small-scale point clouds and not suitable for filtering highly dense and large-scale point clouds. Furthermore, previous works only have demonstrated satisfactory performance solely in low-noise environments, while limited to handling details in high-noise environments. This limitation is particularly evident in the generation of unrealistic boundaries, as well as a lack of appropriate constraints. Consequently, the main challenges are twofold: (i) denoising for highly dense and large-scale point clouds with complex geometries, (ii) denoising for high-noise environments.
@@
Recently, State Space Models (SSMs), particularly structured SSMs (S4) (Gu, Goel, and Ré 2021), have shown great promise in Natural Language Processing (NLP) (Lu et al. 2024). Built upon the S4 framework, Mamba (Gu and Dao 2023) exhibits outstanding performance across many tasks in NLP and 2D CV  (He et al. 2024; Yang et al. 2024; Zhao et al. 2024; Misra, Gala, and Orvieto 2024). The remarkable achievements of Mamba are a testament to its linear computational complexity and exceptional long-range context learning capabilities on extensive sequences. Moreover, the integration of time-varying parameters and a hardware-aware algorithm within Mamba marks a significant leap forward in the field of sequence modeling.
@@
However, exploiting SSM models for point cloud filtering has yet to be fully explored. Direct application of Mamba may not fully address the challenges involved. Therefore, we propose 3DMambaIPF, a novel Iterative Point Clouds Filtering model utilizing the Mamba module with differentiable point rendering techniques. Specifically, 3DMambaIPF is composed of multiple iterations of Mamba-Denoising Modules, each comprising a pair of Mamba encoders and decoders. Initially, the input point cloud is divided into patches, within which a graph structure is generated based on points and their neighbors to extract position features. In the Mamba-based Encoder, the Dynamic EdgeConv module encodes positional features with MLPs, which are then fed into a Mamba module to select and generate position-dependent features from sequential inputs. The Mamba-based Decoder is responsible for upsampling features into point cloud patches. During the training process, a novel rendering loss is introduced leveraging a differential point rendering technique, and the loss is computed by the disparities of rendered images of filtered point cloud and GT. Subsequently, an adaptive GT altering throughout the iterations is introduced, gradually transitioning from an initially noise-added state to the true GT. Following the completion of all iterations, a patch-stitching method is employed to reconstruct the filtered point cloud from the patches. Experimentally, 3DMambaIPF is trained using PU-Net (Yu et al. 2018) dataset and achieves state-of-the-art results compared to off-the-shelf methods. 3DMabaIPF is also able to well generalize to high-noise point clouds from the Stanford 3D Scanning Repository (Curless and Levoy 1996; Krishnamurthy and Levoy 1996) for synthetic objects with complex geometries. Moreover, 3DMambaIPF outperforms baseline methods in terms of visual perceptions on real-world scene datasets. In summary, our main contributions are as follows:
  • A novel iterative Mamba-based backbone is designed for point cloud filtering, allowing for the effective modeling of long-sequence point cloud features. This approach enhances the accuracy and speed of highly dense and large-scale point cloud processing.
@@
  • A well-designed differentiable point rendering loss is introduced into 3DMambaIPF, supplanting the distance-based loss functions. This differentiable rendering loss endows 3DMambaIPF with the ability to handle edge denoising and makes the denoising results more realistic.
@@
  • 3DMambaIPF surpasses existing methods by achieving state-of-the-art results on the commonly used PU-Net dataset, while demonstrating excellent performance even on highly-dense and large-scale datasets with high noise added.
Related Work
@@
Figure 1: Overview of 3DMambaIPF. An Encoder-Decoder-based model built on Mamba named Mamba-Denoising Module is introduced for iterative filtering. The input noisy point cloud is partitioned into patches and fed into the iterative Mamba-Denoising Module. Upon completion of the iterations, clean point cloud patches are produced as output. To enhance the filtering of noisy points around the surface, a differentiable rendering method is introduced. The rendering loss and reconstruction loss are jointly backpropagated to update the parameters of 3DMambaIPF.
@@
Deep Learning-based Point Cloud Filtering Methods. Deep Learning-based point cloud filtering methods are undergoing significant development to address the limitations of traditional approaches. The well-known PCN (PointCleanNet) (Rakotosaona et al. 2020) model is based on PCPNet (Guerrero et al. 2018) and focuses on removing outliers and reducing noise in unordered point clouds. However, due to PCN’s denoising being performed point-wise rather than patch-wise, its efficiency is relatively low. PD-Flow (Mao et al. 2022) learns the distribution of noisy point sets and achieves denoising through noise disentanglement. However, the number of iterations needed for PD-Flow varies depending on the level of noise to achieve optimal results. GPDNet (Pistilli et al. 2020) employs graph-convolutional layers to dynamically build neighborhood graphs and utilizes the similarity of high-dimensional feature representations to construct intricate feature hierarchies. ScoreDenoise (Luo and Hu 2021) leverages the distribution model and score estimation from noisy point clouds. Pointfilter (Zhang et al. 2020) employs an encoder-decoder architecture, where points and their neighbors are taken as input for point-wise learning. It generates a displacement vector to accomplish denoising while preserving sharp features. IterativePFN (de Silva Edirimuni et al. 2023) leverages iterative filtering modules and adaptive GT progressively yet swiftly converges noisy points onto clean surfaces, achieving superior performance. Current point cloud filtering methods are primarily designed for sparse and small-scale datasets (up to 50K points) and struggle to effectively filter dense point clouds (approximately 500K points) with noise far from clean surfaces, particularly in large-scale point clouds. Furthermore, existing methods (Zhang et al. 2020) demonstrate limited effectiveness in filtering the edges of point clouds and capturing fine-grained characteristics. Considering these challenges, our work aims to introduce a model tailored for highly dense and large-scale point clouds characterized by high-density noise and robust filtering capabilities at point cloud edges.
@@
SSM-based Methods. SSMs provide a robust mathematical framework rooted in control theory and are increasingly applied in fields such as NLP and CV to model complex temporal dynamics within dynamic systems. This novel modeling approach addresses the challenge of handling long sequences, a difficulty encountered by both Transformers and RNNs. By mitigating the quadratic computational complexity of Transformers and addressing the forgetting issue of RNNs, SSMs prove highly efficient in capturing temporal dependencies within sequential data.
@@
Significant advancements have been made in the Mamba framework (Gu and Dao 2023), which integrates selective SSMs into a simplified architecture. This integration allows Mamba to achieve rapid inference, linear scalability in sequence length, and state-of-the-art performance across multiple modalities, while also enhancing the selectivity of state representation. These advancements collectively demonstrate the ongoing evolution of SSMs and their increasing significance in modeling temporal dynamics across diverse domains. Methods based on Mamba have achieved state-of-the-art results in various fields, including large language models (He et al. 2024), medical image analysis (Liao et al. 2024), and 3D point cloud perception (Liang et al. 2024). Taking Consideration of the advantages of Mamba in modeling long sequences, we introduce Mamba in our work to address highly-dense and large-scale point cloud modeling problems.
@@
Differentiable Rendering Techniques. Rendering techniques encompass algorithms employed to transform three-dimensional models into two-dimensional images. The differentiable rendering methods are designed for introduction into deep learning, facilitating 3D reconstruction through error backpropagation. Current differentiable rendering methods can be categorized into four types based on their geometric representations: point-based (Yifan et al. 2019; Müller, Weinmann, and Klein 2022; Kerbl et al. 2023; Fei et al. 2024a), triangle mesh-based (Laine et al. 2020; Liu et al. 2019), implicit function-based(Bangaru et al. 2022; Vicini, Speierer, and Jakob 2022), and volume-based(Mildenhall et al. 2021; Ren et al. 2023). In this paper, we propose an efficient differentiable rendering method based on point-based rendering (Insafutdinov and Dosovitskiy 2018) to render both the filtered and GT point clouds during training to generate the view loss, which is then propagated backward to adjust the network parameters.
Method
Overview
Figure 2: A Gaussian noise with a variable standard deviation is added to the GT to generate the adaptive GT. As the standard deviation gradually decreases, the adaptive GT approaches the GT with each iteration. Eventually, the GT no longer changes in the final iteration.
@@
Figure 3: Mamba-Denoising Module for iterative point cloud filtering of 3DMambaIPF. The Mamba-based Encoder comprises Dynamic EdgeConv modules and Mamba Blocks, while the Mamba-based Decoder consists of Mamba Decoder Blocks and an Activated Linear module.
Figure 4: Detailed network architectures for Dynamic EdgeConv, Mamba Block, and Mamba Decoder Block.
@@
The overview of 3DMambaIPF is depicted in Figure 1, which utilizes a novel iterative Mamba-Denosing backbone and a differentiable point rendering technique. Mamba-Denoising backbone consists of a cascade of Mamba-based encoders and decoders, with each pair referred to as a Mamba-Denoising Module. After receiving an extracted patch of the point cloud as input, a single Mamba-Denoising Module constructs a directed graph from the points within the patch. The directed graph is encoded into token sequences to extract features. The denoised point token sequences are then decoded, eliminating noise, and serve as the input for the next Mamba-Denoising Module.
@@
Specifically, given a 3-dimensional clean point cloud, denoted as $X_{c}=\{x_{c1},x_{c2},...x_{cn}\}\subseteq\mathbb{R}^{3}$, each point in $X_{c}$ can be represented by 3-dimensional coordinates $x_{ci}=(\alpha_{i},\beta_{i},\gamma_{i})$. To prepare the point cloud as input for 3DMambaIPF, the K-nearest neighbors (KNN) algorithm (Peterson 2009) is used to partition the point cloud into patches, and a directed graph is constructed within each patch. Considering a point cloud patch $X\subseteq\mathbb{K}^{3}$ which generated by KNN and a directed graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}=\{x_{1},x_{2},...x_{K}\}$ and $\mathcal{E}$ represents the edges between $x_{i}\in\mathcal{V}$ and its $k$-nearest neighbors.
@@
Subsequently, a differentiable 3D point rendering loss is employed to refine the positions of noisy points, especially around the surface of the point clouds, by comparing the rendered images from filtered point clouds with the adaptive GT.
Mamba-Denoising Backbone
@@
The Mamba-Denoising Backbone is the core component of 3DMambaIPF for iterative point cloud filtering. As illustrated in Figure 3, the Mamba-Denoising backbone adopts an Encoder-Decoder structure. The encoder consists of four Dynamic EdgeConv modules and four corresponding Mamba Blocks respectively, while the decoder comprises a Mamba Decoder Block and a linear activation layer. The detailed network structure is illustrated in Figure 4.
@@
The Dynamic EdgeConv module facilitates the conversion of input patches into a graph structure, subsequently processed through an MLP network to extract pertinent features. Initially, vertices undergo feature extraction via an MLP, capturing intrinsic characteristics. Subsequently, an additional MLP focuses on extracting positional features from vertices connected by edges to the target vertex, facilitating feature differentiation. Specifically, vertex features are updated as ${h_{i}}^{l+1}=f({h_{i}}^{l})+\sum_{i,j}g(\text{concat}({h_{i}}^{l},{h_{j}}^{l})-{h_{i}}^{l}),$ where $h^{l}$ denotes to features at layer $l$, $f(\cdot)$, $g(\cdot)$ denote one-layer or two-layers MLPs, $i$ is a vertex on the graph, $(i,j)$ forms an edge. This process culminates in concatenating the target vertex’s features with the differences of features computed from all connected vertices, effectively deriving nuanced positional features within the patch. Following the feature extraction of the point cloud within the patch using MLPs, the resultant feature sequence is sent to the Mamba Encoder. This encoder selectively processes the input feature sequence and alters it into an output feature sequence. In the Decoder module, the Mamba Decoder Block selectively processes the input feature sequence to reduce dimensionality, gradually reconstructing it into three dimensions. Linear activation is used to determine the output movement distance of a noisy point towards a clean surface after filtering. Utilizing a residual connection, the noisy points within the input patch are adjusted under the guidance of the movement distance, resulting in a denoised patch.
Differentiable Rendering and Backpropagation